Concentration (M) of starting solution % Transmittance, %T, (Blank) % T (15-mL 0.024 M Cu2+ + 7 mL conc NH3 diluted with distAbsorbance (15-mL Dilution) NG 2 Info Ratio M/A (use 0.024 M as the given M) for the 15-mL Dilution NG 2 Info Absorbance of UNEED ANSWERS ASAP
Answer #1

@ Absorbance (15 mL Dilution) A=2-log(+2) = 2-log (61.6) = 0.21042 Ablank = 2-log (9:1) = 3193x10-8 Absorbance = 0,21042-3193